Question 20 1 pts When people spend more years in school, this results in a externality because there are created by a more hQuestion 15 1 pts A producer knows that the price elasticity for his product is -0.5. He wants to increase quantity demandedQuestion 24 1 pts Output Fixed (Q) Average Total Marginal Costs Variable Total Average Fixed Average Costs Costs Costs CostsQuestion 25 1 pts Average Total Marginal Output Fixed Variable Total Average Fixed Average Costs (Q) Costs Costs Costs Costs
Answer #1

20. Ans: Positive ; external benefits

Explanation:

Because more educated people tend to bring more economic value to their society.

15. Ans: decrease by 60%

Explanation:

PED = % change in quantity demanded / % change in price

% change in price = % change in quantity demanded / PED

                              = 30% / -0.5

                              = -60%

It means price should decrease by 60%

24.Ans: $700

Explanation:

TC of producing 1 unit of good = FC + VC = $500 + $200 = $700

ATC of producing 1 unit of good = TC / Q = $700 / 1 = $700

25. Ans: 4

Explanation:

Q FC($) VC($) TC($)=FC+VC AFC($)=FC/Q AVC($)=VC/Q ATC($)=TC/Q MC($)=∆TC/∆Q
1 500 200 700 500.00 200.00 700.00
2 500 300 800 250.00 150.00 400.00 100.00
3 500 375 875 166.67 125.00 291.67 75.00
4 500 425 925 125.00 106.25 231.25 50.00
5 500 500 1000 100.00 100.00 200.00 75.00
6 500 625 1125 83.33 104.17 187.50 125.00